OpenStudy (anonymous):

WILL FAN AND MEDAL! a line has a slope of -4 and passes through the point (8,-7). what is the equation of the line?

OpenStudy (anonymous):

(y+7) = -4(x-8)

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (anonymous):

these are the options....

OpenStudy (anonymous):

@pooja195 @maddi07

OpenStudy (anonymous):

@Directrix could you help me plese?:)

OpenStudy (anonymous):

@ArifL do you know which one it is? i posted a picture with four choices on it..

OpenStudy (anonymous):

it the last one

OpenStudy (anonymous):

thank you @ArifL :)